Suppose that Z is the standard normal distribution. Find P(Z<-1.81).

Suppose that Z is the standard normal distribution. Find P(Z>2).

Suppose that Z is the standard normal distribution. Find P(-1.95<Z<1.07).

Suppose that Z is the standard normal distribution. What value of Z represents the 20th percentile?
Answer #1

Solution :

P(Z < -1.81) = 0.0351

P(Z > 2) = 1 - P(Z < 2) = 1 - 0.9772 = 0.0228

P(-1.95 < Z < 1.07) = P(Z < 1.07) - P(Z < -1.95) = 0.8321

P(Z < -0.84) = 0.20

20th percentile = -0.84